Your neighbor Bob has two annuities. The first annuity will pay him $10,000 per month for the next 10 years. The second annuity will pay him $15,000 per month for the following 10 years (years 11 through 20). Assuming a discount rate of 6%, what is the present value of the annuities?

a) $2,251,837

b) $906,288

c) $900,735

d) $3,000,000

e) $1,643,345
Answer #1

e) $1,643,345

Present value of annuity of 1 for 10 years = (1-(1+i)^-n)/i Where,
= (1-(1+0.005)^-120)/0.005 i = 6%/12 =                    0.005
= 90.0734533 n = 10*12 = 120
Present value of 1 of 10 years = (1+i)^-n
= (1+0.005)^-120
= 0.54963273
Present value of cash flows of first 10 years = $ 10,000.00 * 90.07345 = $          9,00,735
Present value of cash flows of next 10 years = $ 15,000.00 * 90.07345 * 0.549633 = $          7,42,610
Present value of cash flows of next 20 years $       16,43,344